OpenStudy (anonymous):

The sum of one fifth of a number and three is equal to half of the number. what number is it?

OpenStudy (anonymous):

basically that means that 1/5*x + 3= 1/2*x so 3=1/2x - 1/5x 3= 0.3x so x=10 :)